The number of vacancies in some hypothetical metal increases by a factor of 4 when the temperature is increased from 900 ˚C to 1130 ˚C. Calculate the energy for vacancy formation (in J/mol) assuming that the density of the metal remains the same over this temperature range.
Answer #1

we know NV/N = exp(-Q/RT), Q = energy for vacancy,

at two different temperatures 900 (T1) and 1130 (T)

(NV/N)2/ (NV/N)1 = exp(- Q/RT2)/ exp(-Q/RT1)

the LHS is fraction by which vacancies increase

4= exp(Q/R)*(1/T1-1/T2)

4= exp(Q/R)*(1/1173-1/1403)

Q = (ln4)*8.314/ (1/1173-1/1403)

Q= 82469 J/mole =82.469 Kj/mole